OpenStudy (anonymous):

Oil is dripping onto a surface at the rate of 1 10cm3/s and forms a circular lm which may be considered to have a uniform depth of 0.1 cm. Find the rate at which the radius of the circular lm is increasing when this radius is 5 cm.

OpenStudy (amistre64):

determine what needs to be solved; and how it relates to the given information

OpenStudy (amistre64):

what is a lm?

OpenStudy (anonymous):

I undertand the problem, but not quite sure what a "circular lm" is... General way to do this problem is to get the formula for the volume of the shape (applying the uniform depth) which will give you a two-variable function. THen take the derivative and for radius = 5.

OpenStudy (anonymous):

edit: "Then take the derivative and solve for when the radius = 5"

OpenStudy (amistre64):

dr/dt is what to find dV/dt is given as 10 V = pi r^2 *.1; derive wrt to time (t) dV/dt = 2pi r * .1 dr/dt 10 = 2pi r * .1 dr/dt \[\frac{10}{2pi\ r(.1)}=\frac{dr}{dt}\]

OpenStudy (anonymous):

lm here is like a big drop... like a pit

OpenStudy (amistre64):

when r=5; solve for dr/dt

OpenStudy (amistre64):

and i think it reads 110 not 10 ...

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

lm was meant to film

OpenStudy (amistre64):

Another way we could have looked at it was from the chain rule: \[\frac{dr}{dt}=\frac{dr}{dV}\frac{dV}{dt}\] \[V = pi\ r^2(.1)\text{ ; derive with respect to V}\] \[1 = 2pi\ r(.1)*\frac{dr}{dV}\] \[\frac{dr}{dV}=\frac{1}{.2pi\ r}\] \[\frac{dr}{dt}=\frac{1}{.2pi\ r}\frac{110}{1}\] same results